# Extremal behavior of stationary marked point processes

### v2

@@ 5.2 Moving maxima model
Let $\zeta$ be a positive random variable whose tail is regularly varying of index $\alpha>0$, and denote the distribution of $\zeta$ by $m$. Assume that $P\in\mathcal{N}_{10}$ is a unit rate independently marked stationary Poisson process on $\mathbb{R}^{d}$ with marks in $(0,\infty)$, whose intensity measure is the product of the Lebesgue measure on $\mathbb{R}^{d}$ and a probability measure $m$, so that $\zeta$ is the typical mark. In the sequel, write $P=\textstyle\sum\delta_{(t,\zeta_{t})}$ where $(\zeta_{t})_{t}$, $t\in\mathbb{R}^{d}$, are i.i.d. with distribution $m$.
Let $\mathcal{N}_{g}$ be the space of all simple locally finite point measures on $\mathbb{R}^{d}$ equipped with the usual $\sigma$-algebra. Consider a measurable function $\Phi:\mathbb{R}^{d}\times\mathcal{N}_{g}\to\mathcal{N}_{g}$ such that, for each $\mu^{\prime}\in\mathcal{N}_{g}$, and each $t\in\mu^{\prime}$, $\Phi(t,\mu^{\prime})$ is a finite subset of $\mu^{\prime}$ which contains $t$. It is useful to interpret $\Phi(t,\mu^{\prime})$ as a neighborhood of $t$ in $\mu^{\prime}$, and all points $x\in\Phi(t,\mu^{\prime})$ as neighbors of $t$. Furthermore, assume that $\Phi$ is shift-equivariant in the sense that $\Phi(t-x,\varphi_{x}\mu^{\prime})=\Phi(t,\mu^{\prime})-x$ for all $\mu^{\prime}\in\mathcal{N}_{g}$, $t\in\mu^{\prime}$ and $x\in\mathbb{R}^{d}$. Finally, denote by $N(t,\mu^{\prime})$ the cardinality of $\Phi(t,\mu^{\prime})$. It can been easily seen that the arguments below also work if $\Phi$ depends on some external sources of randomness, but stays independent of the marks $(\zeta_{t})$.
@@
where $\mu^{\prime}$ is $\mu$ projected to $\mathbb{R}^{d}$. Without loss of generality, in the sequel assume that $w(0)=1$. If the weight function is identically one, the score at $t$ is the maximum of the scores of its neighbors.
Since $\psi$ is shift-invariant, $X:=\Psi(P)$ is a stationary marked point process on $\mathbb{R}^{d}$. Its Palm version is given by $\tilde{X}:=\Psi(\tilde{P})$, where $\tilde{P}=P+\delta_{(0,\zeta_{0})}$, with $\zeta_{0}$ having distribution $m$ and being independent of $P$. For notational convenience, denote $\widetilde{\Phi}(t)=\Phi(t,\tilde{P}^{\prime})$, and let $\tilde{N}(t)=N(t,\tilde{P}^{\prime})$ be the cardinality of $\widetilde{\Phi}(t)$ for all $t\in\tilde{P}^{\prime}$.
@@
for some $a,\varepsilon>0$. We extend $f$ to a continuous function on $\mathbb{R}^{d}\times[0,\infty)$ by letting $f(t,0)=0$ for all $t$. For notational convenience set $h(\mu):=e^{-\mu(f)}$ for all $\mu\in\mathcal{N}_{11}$. The result would follow if we show that
@@
First, write $\tilde{P}=\{(t_{i},\zeta_{i}):i\geq 1\}$, where $\{t_{1},\dots,t_{\tilde{N}(0)}\}=\widetilde{\Phi}(0)$. Given $\tilde{P}^{\prime}$, the score $\psi(t,\tilde{P})$ depends only on $\zeta_{x}$ for $x\in\widetilde{\Phi}(t)$. By (5.8),
@@
$g_{u}(\mu)=\mathbf{E}\Big{[}g\big{(}\mu^{\prime},\zeta_{1}/u,\dots,\zeta_{K_{a}(\mu^{\prime})}/u\big{)}\boldsymbol{1}{\left\{\textstyle\max_{i\leq N(0,\mu^{\prime})}w(t_{i})\zeta_{i}>u\right\}}\Big{]},$ (5.12)
@@
Since $(\zeta_{i})_{i}$ are i.i.d. regularly varying with index $\alpha$, the vector $(\zeta_{1},\dots,\zeta_{k})$ is multivariate regularly varying in +$\mathbb{R}^{k}_{+}\setminus\{0\}$ with the same index, see [24, p. 192]. In particular,
@@
for a certain measure $\nu$ on $\mathbb{R}_{+}^{k}\setminus\{0\}$ concentrated on the axes. More precisely, by [24, p. 192], if $\eta$ is $\mathrm{Pareto}(\alpha)$ distributed, the right-hand side of (5.13) equals
@@
almost surely as $u\to\infty$. Moreover, since $\mathbf{E}[\tilde{N}(0)w_{*}^{\alpha}]=\mathbf{E}[\tilde{N}(0)]w_{*}^{\alpha}$ is finite by assumption, Pratt’s extension of the dominated convergence theorem (see [23, Theorem 1]) justifies the interchange
@@
In the following assume that the assumptions of Theorem 5.4 hold so that $X$ admits a tail configuration $Y$ whose spectral configuration $\Theta$ is given by (5.5); recall the convention explained in Remark 5.5. In order to determine the ingredients of the limiting point process in Theorem 4.2, the following result is crucial; our approach used here is similar to the one taken by [22, Section 4] for random fields over $\mathbb{Z}^{d}$.
@@
We now give sufficient conditions under which the assumptions of Theorem 4.2 are satisfied. First, take a family $(a_{\tau})_{\tau}$ such that (4.1) holds, and fix an arbitrary family $(b_{\tau})_{\tau}$ such that $b_{\tau}\to\infty$ and $b_{\tau}/\tau\to 0$ as $\tau\to\infty$; since $r$ is a constant function, this is equivalent to choosing $(b_{\tau})$ such that 4.3 holds.
@@
    Assume that for some $r_{0}>0$, with probability $1$,
    $\Phi(t,\tilde{P}^{\prime})=\Phi(t,\tilde{P}^{\prime}\cap B(t,r_{0}))\,,\;\;\text{for all $t\in\tilde{P}^{\prime}$}\,.$
@@
    If $\Phi(t,\mu^{\prime})$ is the set containing $t$ and the $k$ nearest neighbors of $t$ in $\mu^{\prime}\setminus\{t\}$ (with respect to the Euclidean distance), $R$ satisfying 5.10 can be constructed as in the proof of [20, Lemma 6.1], see also [12, p. 104]. Observe that in this case taking $R(\mu^{\prime})$ to be the distance to the $k$-th nearest neighbor of 00 in $\mu^{\prime}\setminus\{0\}$ is not sufficient for property (i), and this property is crucial to ensure that the anticlustering condition holds.
@@
If 5.10 holds, 4.5 is satisfied for any $(l_{\tau})_{\tau}$ such that $l_{\tau}\to\infty$ and $l_{\tau}/b_{\tau}\to 0$ as $\tau\to\infty$.
@@
for all $\tau>0$, $\boldsymbol{i}\in I_{\tau}$. To confirm 4.5, we need to show that
@@
Assume that $(b_{\tau})_{\tau>0}$ satisfies Assumptions 4.3 and 4.4. Then
@@
Since $b_{\tau}/r(a_{\tau})\to\infty$ by 4.3, the two previous relations imply that
@@
Let $u,\varepsilon>0$ be arbitrary. Due to the first inclusion in (6.5) and since $b_{\tau}/r(a_{\tau}y)\to\infty$ by 4.3, $B_{r(a_{\tau}y)u}\subseteq D_{\tau}$ for all sufficiently large $\tau$. Observe that if
@@
where the last inequality holds, since $D_{\tau}\subseteq B_{b_{\tau}c_{2}}$. If now one lets $u\to\infty$ and then $\varepsilon\to 0$, 4.4 implies (6.13), and this proves (6.4). ∎
@@
In particular, $A_{1}$ is precisely the first exceedance anchoring function $A^{\mathrm{fe}}$ from (4.20). As shown in Lemma 4.6, $\mathbf{P}(A_{1}(Y)=0)$ is positive whenever $\mathbf{P}(Y\in\mathcal{N}_{01})=1$, which holds, for instance, under Assumptions 4.3 and 4.4. Recall also that $X_{\tau}:=X_{[0,b_{\tau}]^{d}}$.
@@
Assume that Assumptions 4.3 and 4.4 hold. Then for every $y>0$,
@@
Observe here that $f$ is not continuous on the whole $\mathcal{N}_{01}$, but since the probability that $(t,1)\in Y$ for some $t\in\mathbb{R}^{d}$ is zero (due to Propositions 3.2 and 3.3), it is continuous on the support of $Y$ – this justifies the use of (6.4). By the dominated convergence theorem and since the limit in (6.22) does not depend on $s$, (6.21) yields that
@@
Assume that Assumptions 4.3 and 4.4 hold. Then
@@
We now prove (6.30). By [14, Lemma 4.1], it suffices to prove that $\mathbf{E}\big{[}N_{\tau}(g)\big{]}\to(\operatorname{Leb}\times\nu)(g)$ for all $g:[0,1]^{d}\times\tilde{\mathcal{N}}_{01}^{*}\to[0,\infty)$ of the form $g(t,[\mu])=\boldsymbol{1}_{(a,b]}(t)\boldsymbol{1}_{A}([\mu])$, where $(a,b]$ is the parallelepiped in $\mathbb{R}^{d}$ determined by $a=(a_{1},\dots,a_{d}),b=(b_{1},\dots,b_{d})\in[0,1]^{d}$ with $a_{j}\leq b_{j}$ for all $j$, and $A\in\mathcal{B}^{*}$ (with $\mathcal{B}^{*}$ defined just before Proposition 6.3) such that $\nu(\partial A)=0$.
@@
Assume that $(b_{\tau})_{\tau>0}$ is such that Assumptions 4.3 and 4.4 hold. Then, Assumption 4.5 implies
@@
By the construction of $f^{\delta}$ and properties of $f$, 4.5 implies that (4.19) holds for the family
@@
Since the family of Lipschitz continuous functions from Lemma 6.4 is convergence determining in the sense of [3, Definition 2.1] (see [2, Proposition 4.1]), the convergence of intensities (6.30) and the asymptotic independence of blocks (6.33) imply that, as $\tau\to\infty$, $N_{\tau}$ converges in distribution to a Poisson point process $N$ on $\boldsymbol{N}=[0,1]^{d}\times\tilde{\mathcal{N}}_{01}^{*}$ whose intensity measure is $\operatorname{Leb}\times\nu$; this is [3, Theorem 2.1] which is a consequence of the classical Grigelionis theorem, see [14, Corollary 4.25]. Standard transformation results for Poisson processes now imply that $N$ can be constructed as in (4.11), and this finishes the proof of Theorem 4.2. ∎
@@ 7 Acknowledgement
This work was supported by the Swiss Enlargement Contribution in the framework of the Croatian–Swiss Research Programme (project number IZHRZ0_180549).

### v3

@@ 5.2 Moving maxima model
Moving maxima models are very popular in the studies of time series. Below we put such models in the spatial context by starting with a marked Poisson point process and then adjusting the scores by giving each point the score derived from weighted maximum of the scores of points from its neighbours. This general construction was described in Section 2.2.
Let $\zeta$ be a positive random variable whose tail is regularly varying of index $\alpha>0$, and denote the distribution of $\zeta$ by $m$. Assume that $P\in\mathcal{N}_{10}$ is a unit rate independently marked stationary Poisson process on $\mathbb{R}^{d}$ with marks in $(0,\infty)$, whose intensity measure is the product of the Lebesgue measure on $\mathbb{R}^{d}$ and a probability measure $m$, so that $\zeta$ is the typical mark. In the sequel, write $P=\textstyle\sum\delta_{(t,\zeta_{t})}$ where $(\zeta_{t})_{t}$, $t\in\mathbb{R}^{d}$, are i.i.d. with distribution $m$. While, by independence, the large values of $\zeta_{t}$ come in isolation in $P$, the clustering of scores is easily modeled by considering, for instance, $X=\textstyle\sum\delta_{(t,\max_{|t-s|<1}\zeta_{s})}$, where the “large” scores in $P$ propagate to all neighbouring locations in the ball of radius 1.
Let $\mathcal{N}_{g}$ be the space of all simple locally finite point measures on $\mathbb{R}^{d}$ equipped with the usual $\sigma$-algebra. Consider a measurable function $\Phi:\mathbb{R}^{d}\times\mathcal{N}_{g}\to\mathcal{N}_{g}$ such that, for each $\mu^{\prime}\in\mathcal{N}_{g}$, and each $t\in\mu^{\prime}$, $\Phi(t,\mu^{\prime})$ is a finite subset of $\mu^{\prime}$ which contains $t$. It is useful to interpret $\Phi(t,\mu^{\prime})$ as a neighborhood of $t$ in $\mu^{\prime}$, and all points $x\in\Phi(t,\mu^{\prime})$ as neighbours of $t$. Furthermore, assume that $\Phi$ is shift-equivariant in the sense that $\Phi(t-x,\varphi_{x}\mu^{\prime})=\Phi(t,\mu^{\prime})-x$ for all $\mu^{\prime}\in\mathcal{N}_{g}$, $t\in\mu^{\prime}$ and $x\in\mathbb{R}^{d}$. Finally, denote by $N(t,\mu^{\prime})$ the cardinality of $\Phi(t,\mu^{\prime})$. It can been easily seen that the arguments below also work if $\Phi$ depends on some external sources of randomness, but stays independent of the marks $(\zeta_{t})$.
@@
where $\mu^{\prime}$ is the projection of $\mu$ on $\mathbb{R}^{d}$. Without loss of generality, in the sequel assume that $w(0)=1$. If the weight function is identically one, the score at $t$ is the maximum of the scores of its neighbours.
Since $\psi$ is shift-invariant, $X:=\Psi(P)$ is a stationary marked point process on $\mathbb{R}^{d}$. Its Palm version is given by $\tilde{X}:=\Psi(\tilde{P})$, where $\tilde{P}=P+\delta_{(0,\zeta_{0})}$, with $\zeta_{0}$ having distribution $m$ and being independent of $P$. For notational convenience, denote $\widetilde{\Phi}(t)=\Phi(t,\tilde{P}^{\prime})$, and let $\tilde{N}(t)=N(t,\tilde{P}^{\prime})$ be the cardinality of $\widetilde{\Phi}(t)$ for all $t\in\tilde{P}^{\prime}$. Clearly, $X=\textstyle\sum\delta_{(t,\max_{|t-s|<1}\zeta_{s})}$ is a special example of this construction with $w\equiv 1$ and $\Phi(t,\mu^{\prime})=\textstyle\sum_{s\in\mu^{\prime}}\delta_{s}\boldsymbol{1}_{\{|s-t|<1\}}$, i.e., for the neighbourhoods which consist of all points $s\in\mu^{\prime}$ within distance less than $1$ to $t$.
@@
Recall that the random score $\zeta$ is assumed to be regularly varying with tail index $\alpha$.
@@
for some $a,\varepsilon>0$. We extend $f$ to a continuous function on $\mathbb{R}^{d}\times[0,\infty)$ by letting $f(t,0)=0$ for all $t$. For notational convenience set $h(\mu):=e^{-\mu(f)}$ for all $\mu\in\mathcal{N}_{11}$. Both (i) and (ii) would follow immediately if we show that
@@
First, write $\tilde{P}=\{(t_{i},\zeta_{i}):i\geq 1\}$, where $\{t_{1},\dots,t_{\tilde{N}(0)}\}=\widetilde{\Phi}(0)$. Given the projection $\tilde{P}^{\prime}$ of $\tilde{P}$, the score $\psi(t,\tilde{P})$ depends only on $\zeta_{x}$ for $x\in\widetilde{\Phi}(t)$. By (5.8),
@@
$g_{u}(\mu^{\prime}):=\mathbf{E}\Big{[}g\big{(}\mu^{\prime},\zeta_{1}/u,\dots,\zeta_{K_{a}(\mu^{\prime})}/u\big{)}\boldsymbol{1}{\left\{\textstyle\max_{i\leq N(0,\mu^{\prime})}w(t_{i})\zeta_{i}>u\right\}}\Big{]},$ (5.12)
@@
Since $(\zeta_{i})_{i}$ are i.i.d. regularly varying with index $\alpha$, the vector $(\zeta_{1},\dots,\zeta_{k})$ is multivariate regularly varying in +$\mathbb{R}^{k}_{+}\setminus\{0\}$ with the same index, see [26, p. 192]. In particular,
@@
for a certain measure $\nu$ on $\mathbb{R}_{+}^{k}\setminus\{0\}$ concentrated on the axes. More precisely, by [26, p. 192], if $\eta$ is $\mathrm{Pareto}(\alpha)$ distributed, the right-hand side of (5.13) equals
@@
almost surely as $u\to\infty$. Moreover, since $\mathbf{E}[\tilde{N}(0)w_{*}^{\alpha}]=\mathbf{E}[\tilde{N}(0)]w_{*}^{\alpha}$ is finite by assumption, Pratt’s extension of the dominated convergence theorem (see [25, Theorem 1]) justifies the interchange
@@
In the following assume that the assumptions of Theorem 5.4 hold so that $X$ admits a tail configuration $Y$ whose spectral configuration $\Theta$ is given by (5.5); recall the convention explained in Remark 5.5. In order to determine the ingredients of the limiting point process in Theorem 4.2, the following result is crucial; our approach used here is similar to the one taken by [24, Section 4] for random fields over $\mathbb{Z}^{d}$.
@@
We now give sufficient conditions under which the assumptions of Theorem 4.2 are satisfied. First, take a family $(a_{\tau})_{\tau}$ such that (4.1) holds, and fix an arbitrary family $(b_{\tau})_{\tau}$ such that $b_{\tau}\to\infty$ and $b_{\tau}/\tau\to 0$ as $\tau\to\infty$; since $r$ is a constant function, this is equivalent to choosing $(b_{\tau})$ such that 4.4 holds.
@@
    Assume that for some $r_{0}>0$,
    $\Phi(t,\mu^{\prime})=\{x\in\mu^{\prime}:|t-x|<r_{0}\}\,,\;\;\text{for all $t\in\mu^{\prime}$}\,.$
@@
    If $\Phi(t,\mu^{\prime})$ is the set containing $t$ and the $k$ nearest neighbors of $t$ in $\mu^{\prime}\setminus\{t\}$ (with respect to the Euclidean distance), $R$ satisfying 5.10 can be constructed as in the proof of [22, Lemma 6.1], see also [14, p. 104]. Observe that in this case taking $R(\mu^{\prime})$ to be the distance to the $k$-th nearest neighbor of 00 in $\mu^{\prime}\setminus\{0\}$ is not sufficient for property (i), and this property is crucial to ensure that the anticlustering condition holds.
@@
If 5.10 holds, 4.6 is satisfied for any $(l_{\tau})_{\tau}$ such that $l_{\tau}\to\infty$ and $l_{\tau}/b_{\tau}\to 0$ as $\tau\to\infty$.
@@
for all $\tau>0$, $\boldsymbol{i}\in I_{\tau}$. To confirm 4.6, we need to show that
@@
Assume that $(b_{\tau})_{\tau>0}$ satisfies Assumptions 4.4 and 4.5. Then
@@
Since $b_{\tau}/r(a_{\tau})\to\infty$ by 4.4, the two previous relations imply that
@@
Let $u,\varepsilon>0$ be arbitrary. Due to the first inclusion in (6.5) and since $b_{\tau}/r(a_{\tau}y)\to\infty$ by 4.4, $B_{r(a_{\tau}y)u}\subseteq D_{\tau}$ for all sufficiently large $\tau$. Observe that if
@@
where the last inequality holds, since $D_{\tau}\subseteq B_{b_{\tau}c_{2}}$. If now one lets $u\to\infty$ and then $\varepsilon\to 0$, 4.5 implies (6.13), and this proves (6.4). ∎
@@
In particular, $A_{1}$ is precisely the first exceedance anchoring function $A^{\mathrm{fe}}$ from (4.20). As shown in Lemma 4.8, $\mathbf{P}(A_{1}(Y)=0)$ is positive whenever $\mathbf{P}(Y\in\mathcal{N}_{01})=1$, which holds, for instance, under Assumptions 4.4 and 4.5. Recall also that $X_{\tau}:=X_{[0,b_{\tau}]^{d}}$.
@@
Assume that Assumptions 4.4 and 4.5 hold. Then for every $y>0$,
@@
Observe here that $f$ is not continuous on the whole $\mathcal{N}_{01}$, but since the probability that $(t,1)\in Y$ for some $t\in\mathbb{R}^{d}$ is zero (due to Propositions 3.3 and 3.4), it is continuous on the support of $Y$ – this justifies the use of (6.4). By the dominated convergence theorem and since the limit in (6.22) does not depend on $s$, (6.21) yields that
@@
Assume that Assumptions 4.4 and 4.5 hold. Then
@@
We now prove (6.30). By [16, Lemma 4.1], it suffices to prove that $\mathbf{E}\big{[}N_{\tau}(g)\big{]}\to(\operatorname{Leb}\times\nu)(g)$ for all $g:[0,1]^{d}\times\tilde{\mathcal{N}}_{01}^{*}\to[0,\infty)$ of the form $g(t,[\mu])=\boldsymbol{1}_{(a,b]}(t)\boldsymbol{1}_{A}([\mu])$, where $(a,b]$ is the parallelepiped in $\mathbb{R}^{d}$ determined by $a=(a_{1},\dots,a_{d}),b=(b_{1},\dots,b_{d})\in[0,1]^{d}$ with $a_{j}\leq b_{j}$ for all $j$, and $A\in\mathcal{B}^{*}$ (with $\mathcal{B}^{*}$ defined just before Proposition 6.3) such that $\nu(\partial A)=0$.
@@
Assume that $(b_{\tau})_{\tau>0}$ is such that Assumptions 4.4 and 4.5 hold. Then, Assumption 4.6 implies
@@
By the construction of $f^{\delta}$ and properties of $f$, 4.6 implies that (4.19) holds for the family
@@
Since the family of Lipschitz continuous functions from Lemma 6.4 is convergence determining in the sense of [3, Definition 2.1] (see [2, Proposition 4.1]), the convergence of intensities (6.30) and the asymptotic independence of blocks (6.33) imply that, as $\tau\to\infty$, $N_{\tau}$ converges in distribution to a Poisson point process $N$ on $\boldsymbol{N}=[0,1]^{d}\times\tilde{\mathcal{N}}_{01}^{*}$ whose intensity measure is $\operatorname{Leb}\times\nu$; this is [3, Theorem 2.1] which is a consequence of the classical Grigelionis theorem, see [16, Corollary 4.25]. Standard transformation results for Poisson processes now imply that $N$ can be constructed as in (4.11), and this finishes the proof of Theorem 4.2. ∎
@@ 7 Acknowledgement
This work was supported by the Swiss Enlargement Contribution in the framework of the Croatian–Swiss Research Programme (project number IZHRZ0_180549), and by the Croatian science foundation project IP-2022-10-2277.